OpenStudy (anonymous):

One morning, Justin drove 3 hours before stopping to eat lunch. After lunch, he increased his speed by 10 mph. If he completed a 390-mile trip in 6 hours of driving time, how fast (in miles per hour) did he drive in the morning?

OpenStudy (kkutie7):

before lunch justin drove xmiles/hour for 3hours. after lunch he drove x+10miles/hour for 3hours. the addition of this is 390miles/6hours my first frame of thought

OpenStudy (kkutie7):

\[(\frac{xmiles}{hour}*\frac{3hours}{1})+(\frac{x+10miles}{hour}*\frac{3hours}{1})=\frac{390miles}{6hours}\] again just an idea

OpenStudy (anonymous):

it wasnt right but it helped me get the right answer. it would have been correct if you didnt include the 6 hours THANKS !!!

OpenStudy (kkutie7):

oh right =) well at least I helped lol